OpenStudy (anonymous):

x minus eight over three = twelve 28 23 20 44

OpenStudy (anonymous):

\[\frac{x-8}{3}=12\]\[3*\frac{x-8}{3}+8=(12*3)+8\]\[x=44\]

OpenStudy (anonymous):

Can you help me out with one more please?:)

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 3x < -9 or x + 2 10 A number line with a closed circle on -3, shading to the left, and an open circle on 8, shading to the right. A number line with an open circle on -3, shading to the left, and a closed circle on 8, shading to the right. A number line with an open circle on -3, a closed circle on 8, and shading in between. A number line with a closed circle on -3, an open circle on 8, and shading in between.